40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 16 02:27:56 2013, Jenne, Update, LSC, PRMI + 2 arm attempt PRCL_OLTF_PRMIir_XARMgreen_16Oct2013.pdfMICH_OLTF_PRMIir_XARMgreen_16Oct2013.pdfTimeSeries_PRMI_2arms_16Oct2013.png
    Reply  Wed Oct 16 02:39:24 2013, rana, Update, LSC, PRMI + 2 arm attempt 
    Reply  Wed Oct 16 17:34:28 2013, Jenne, Update, LSC, PRMI + 2 arm attempt TimeSeries_PRMI_Xarm_16Oct2013.png
Message ID: 9244     Entry time: Wed Oct 16 02:39:24 2013     In reply to: 9243
Author: rana 
Type: Update 
Category: LSC 
Subject: PRMI + 2 arm attempt 

 

 that's some hot stuff

its time to get the CM servo hardware turned back on. We're going to want to switch it on when we're about ~1/50th of the way up the CARM fringe.

A good way to re-commission it is to lock it to the single arm, using a Pomona box filter to move the arm pole down to the coupled cavity pole frequency.

 

ELOG V3.1.3-